Suppose the total cost function for a firm is given by: TC= 100 + 2q +4q2.

Find the marginal cost function and then use that to determine the marginal cost of the 10th unit.
Answer #1

Marginal cost is the change in total cost due to one unit change in output produced. Mathematically, it s expressed as the following

MC = d(TC)/dq

= 2+ 8q

For q = 10 ,

MC = 2+ 8* 10

=82